OpenStudy (anonymous):

Write an equation of the line that passes through (8, 3) and (4, 3).

jimthompson5910 (jim_thompson5910):

m=(3-3)/(4-8) m=(0)/(4-8) m=(0)/(-4) m=0 -------------------------------------------------------------------------- y-y1 = m(x-x1) y-3=0(x-8) y-3=0x+0(-8) y-3=0x+0 y-3=0 y=0+3 y = 3 So the equation of the line is y = 3

jimthompson5910 (jim_thompson5910):

A quick way to get the answer is to notice that the y coordinates are equal to 3. So the equation is y = 3